Wilma is considering opening a widget factory. The unlevered cost of equity for making widgets is 0.13. This factory would cost $23 million to set up, and would produce EBIT of $3 million per year for the foreseeable future. She is thinking of applying for a $5 million subsidized perpetual loan to finance this project. Complying with the auditing requirements of this loan would have a present value of $2 million. This loan would have a rate of 0.05, while the rate she could get from the bank is 0.07. Her tax rate is 0.31. What is the NPV of this project, using the APV method?
Answer #1
Adjusted Present Value
The adjusted present value is the net present value (NPV) of a project or company if financed solely by equity plus the present value (PV) of any financing benefits, which are the additional effects of debt. By taking into account financing benefits, APV includes tax shields such as those provided by deductible interest.
Adjusted Present Value = Unlevered Firm Value + Net Effect of Debt

Where the net effect of debt includes: Interest tax shield that's created when companies have debt because the interest on the debt is tax-deductible. It’s calculated as interest expense times tax rate. But it only includes the interest tax shield used during that year.
Then the present value of the interest tax shield used must be calculated. The present value of the interest tax shield is: (tax rate * debt load * interest rate) / interest rate.
1) Unlevered Firm Value = An unlevered firm carries no debt and is financed completely through equity.
The equation to calculate the value of an unlevered firm is: [(pre-tax earnings)(1-corporate tax rate)] / the required rate of return.
a) Pre Tax Earnings /(Earnings Before Tax)
EBIT (Earnings before interest and tax) = Rs. 3 Mn per year
Interest = Rs. 5 Mn (Loan) * Rate of Interest 5% = 0.25 Mn
EBT (Earnings Before Tax) = (3 - 0.25) = 2.75 Mn
a) EBT = 2.75
b) Corporate Tax Rate = 31%
c) Required Rate of Return on Equity 13%
d) Unlevered Firm Value         14.60
(d = a * (1 - b )/ c)
2) Net Effect of Debt = NPV of Debt + NPV of Interest Tax Shield
NPV of Debt = Rs. 2 Mn
NPV of Interest Tax Shield = (tax rate * debt load * interest rate) / interest rate.
a) Corporate Tax Rate = 31%
b) Debt Load           5.00 Mn
c) Interest Rate 5%
d) NPV of Interest Tax Shield           1.55 Mn
( d = ( a * b * c) / c
Net Effect of Debt = NPV of Debt + NPV of Interest Tax Shield
Net Effect of Debt = 2 Mn + 1.55 MN = 3.55 Mn
Adjusted Present Value (APV) = Unlevered Firm Value + Net Effect of Debt
Therefore, APV = 14.6 + 3.55 = 18.15 Mn